OpenStudy (vera_ewing):

When looking at a rational function, Charles and Bobby have two different thoughts. Charles says that the function is defined at x = –2, x = 3, and x = 5. Bobby says that the function is undefined at those x values. Describe a situation where Charles is correct, and describe a situation where Bobby is correct. Is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (vera_ewing):

@freckles Do you know how to do this one?

OpenStudy (freckles):

for undefined I think it is easiest to think about vertical asymptotes or holes

OpenStudy (freckles):

Like f(x)=x-3 exists everywhere f(3)=3-3=0 but g(x)=1/(x-3) exists everywhere but at x=3 sincve g(3)=1/0 which is not defined there is a vertical asymptote at x=3

OpenStudy (freckles):

we can also look at hole instead of va if you wish h(x)=(x-3)/(x-3) doesn't exist at x=3 because the bottom would be 0 at x=3 but instead of there being a vertical asymptote at x=3 we have a hole at x=3 since (x-3)/(x-3)=1 when x isn't 3 h(x)=1 when x isn't 3

OpenStudy (vera_ewing):

So now we have to describe a situation where Charles is correct, and describe a situation where Bobby is correct. and then say if it is it possible for a situation to exist where they are both correct? Justify your reasoning.

OpenStudy (freckles):

I think you can now describe a way where charles is correct and also a way where bobby is right but you can't have that that a number is defined and not defined it is either defined or undefined but never both

OpenStudy (vera_ewing):

I'm not sure how to answer the question. Can you help explain what it would be?

OpenStudy (freckles):

can you tell me what you don't understand above in what I wrote?

OpenStudy (freckles):

like can you tell me where this function is undefined? \[f(x)=\frac{1}{(x-4)(x-6)(x+6)}\]

OpenStudy (freckles):

what numbers can you not plug into the bottom?

OpenStudy (vera_ewing):

I just don't get how to describe a situation where Charles is correct, and describe a situation where Bobby is correct.

OpenStudy (vera_ewing):

First, what would be a situation where Charles is correct? Then, what would be a situation where Bobby is correct?

OpenStudy (freckles):

can you answer my question please?

OpenStudy (vera_ewing):

Oh well I guess you can't plug in 0 to the bottom?

OpenStudy (freckles):

well zero is perfectly fine the plug into my function because the bottom will not be zero when replacing x with zero

OpenStudy (freckles):

when is x-4 equal to 0? when is x+6 equal to 0? when is x-6 equal to 0?

OpenStudy (vera_ewing):

Oh gosh I don't know... I'm sorry I'm so confused :(

OpenStudy (freckles):

x-4=0 I know you can solve this equation

OpenStudy (freckles):

even if you can't solve it I bet you should be able to give me a number that when you take 4 from it you get 0

OpenStudy (vera_ewing):

Wait so is that it? Like is that what I would say for the answer?

OpenStudy (freckles):

I don't get what your asking what is it?

OpenStudy (freckles):

\[f(x)=\frac{1}{(x-4)(x+6)(x-6)} \text{ is undefined at } x=4,-6,6 \] because all of these at one point make the bottom 0

OpenStudy (vera_ewing):

How do I answer the question? Can you tell me what I would say for my answer?

OpenStudy (freckles):

do you not see that 4-4=0 and -6+6=0 and 6-6=0?

OpenStudy (vera_ewing):

No I get that, I just need to know what to say for my answer, that's all.

OpenStudy (freckles):

like if x=4 you get f(4)=1/0 which is not defined

OpenStudy (freckles):

so you can now describe a situation where bobby is right ?

OpenStudy (freckles):

just make up a function that is undefined at the values you listed in your question

OpenStudy (vera_ewing):

Can you show me?

OpenStudy (freckles):

I tried above.

OpenStudy (freckles):

I gave you an example

OpenStudy (vera_ewing):

Oh ok sorry I didn't see that. I get it now. Thanks! :)